OpenStudy (anonymous):

In the problems 1 to 4, write the expression |x-2|+|x-5| without the symbols of "Absolute Value", if x are in the intervals: 1. (-∞, 1) 2. (7, ∞) 3. (3, 4] 4. [2, 5]

OpenStudy (anonymous):

can't wait to see this!

myininaya (myininaya):

if x-2>0 or x>2 then |x-2|=x-2 if x-2<0 or x<2 then |x-2|=-(x-2) if x-5>0 or x>5 then |x-5|=x-5 if x-5<0 or x<5 then |x-5|=-(x-5) -(x-2) x-2 x-2 -(x-5) -(x-5) x-5 ---------|-------|-------- 2 5

myininaya (myininaya):

so (-inf,1) we have -(x-2)-(x-5)

myininaya (myininaya):

(7,inf) we have x-2+x-5

myininaya (myininaya):

(3,4] we have x-2-(x-5)

OpenStudy (anonymous):

1. x<1 2. 7<x 3. 3<x<=4 4. 2<=x<=5

myininaya (myininaya):

[2,5] we have x-2-(x-5)

myininaya (myininaya):

wait something is wrong

OpenStudy (anonymous):

no you right, the first one is -2x-7

myininaya (myininaya):

wait i think it is right lol

myininaya (myininaya):

the first one is -2x+7 i think you forgot to distribute the negative 1

myininaya (myininaya):

so (-inf,1) we have -(x-2)-(x-5)=-x+2-x+5=-2x+7

myininaya (myininaya):

(7,inf) we have x-2+x-5=2x-7

OpenStudy (anonymous):

oh yes, thanks!

myininaya (myininaya):

(3,4] we have x-2-(x-5)=x-2-x+5=-2+5=3

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

i cant give you more medals :(

myininaya (myininaya):

lol its k